# CLIPArTT: Light-weight Adaptation of CLIP to New Domains at Test Time

### v1

@@ 5.2 Experiments
Performance analysis under natural or no domain shift.  In the different Tables 8, 9, 10, and 11, CLIPArTT consistently enhances accuracy across CIFAR-10, CIFAR10.1, and CIFAR-100 datasets compared to the baseline (+2%, +3%, +8% respectively with ViT-B/32). This improvement suggests that even datasets devoid of corruption, which may not have been seen by the model during training, can benefit from adaptation. Consequently, a zero-shot model can leverage adaptation at test-time, relying solely on the model’s predictions. However, while CLIPArTT outperforms the baseline, as well as LAME and TPT with this fixed version, it is worth noting that TENT may yield superior results depending on the visual encoder and the number of classes. Entropy minimization used solely on confident results proves to be also effective. Interestingly, the corruptions in which TENT achieves a better performance tend to already have higher accuracy values.
CLIP LAME TENT TPT (BS=32) CLIPArTT
Original 61.68 58.27 $\pm$0.17 69.74 $\pm$0.16 63.78 $\pm$0.28 69.79 $\pm$0.04
Gaussian Noise 14.80 12.72 $\pm$0.04 14.38 $\pm$0.14 14.03 $\pm$0.10 25.32 $\pm$0.14
Shot noise 16.03 13.78 $\pm$0.08 17.34 $\pm$0.27 15.25 $\pm$0.17 27.90 $\pm$0.05
Impulse Noise 13.85 07.82 $\pm$0.14 10.03 $\pm$0.13 13.01 $\pm$0.13 25.62 $\pm$0.09
Defocus blur 36.74 33.38 $\pm$0.11 49.05 $\pm$0.07 37.60 $\pm$0.17 49.88 $\pm$0.23
Glass blur 14.19 09.00 $\pm$0.05 03.71 $\pm$0.07 16.41 $\pm$0.02 27.89 $\pm$0.03
Motion blur 36.14 32.79 $\pm$0.13 46.62 $\pm$0.27 37.52 $\pm$0.23 47.93 $\pm$0.14
Zoom blur 40.24 37.57 $\pm$0.15 51.84 $\pm$0.15 42.99 $\pm$0.11 52.70 $\pm$0.06
Snow 38.95 35.49 $\pm$0.18 46.71 $\pm$0.21 42.35 $\pm$0.13 49.72 $\pm$0.01
Frost 40.56 37.22 $\pm$0.21 44.90 $\pm$0.27 43.31 $\pm$0.14 49.63 $\pm$0.12
Fog 38.00 35.94 $\pm$0.09 47.31 $\pm$0.04 38.81 $\pm$0.17 48.77 $\pm$0.04
Brightness 48.18 44.93 $\pm$0.08 60.58 $\pm$0.18 50.23 $\pm$0.11 61.27 $\pm$0.08
Contrast 29.53 27.52 $\pm$0.06 45.90 $\pm$0.11 28.09 $\pm$0.09 48.55 $\pm$0.24
Elastic transform 26.33 24.01 $\pm$0.02 33.09 $\pm$0.08 28.12 $\pm$0.15 37.45 $\pm$0.08
Pixelate 21.98 19.55 $\pm$0.13 26.47 $\pm$0.09 20.43 $\pm$0.14 33.88 $\pm$0.14
JPEG compression 25.91 21.77 $\pm$0.14 29.89 $\pm$0.07 28.82 $\pm$0.09 36.07 $\pm$0.32
Average 29.43 26.23 35.19 30.46 41.51
Table 9: Accuracy (%) on CIFAR-100 and CIFAR-100-C datasets with ViT-B/32 as visual encoder.
@@
Performance analysis under common corruptions.  In the various referenced tables, including Tables 8, 9, 10, and 11, CLIPArTT consistently outperforms the baseline, TPT, and LAME across all visual encoders, resulting in enhancements of up to 13% on average. A more detailed examination reveals that CLIPArTT frequently surpasses TENT, particularly in instances where the baseline performs poorly. This observation implies that CLIPArTT consistently outperforms TENT, which tends to compromise the model’s performance under conditions of low baseline confidence. For instance, in Table 9, TENT experiences a 3% decrease compared to the baseline under Impulse Noise, while CLIPArTT exhibits a 12% improvement. Conversely, when the baseline is confident, CLIPArTT consistently maintains a high performance. As an illustration, Table 8 shows that while CLIP achieves an accuracy of 81.87% under Brightness corruption, both TENT and CLIPArTT contribute to improvements of 7% and 6%, respectively.
CLIP TENT TPT (BS=32) CLIPArTT
Original ViT-B/16 89.25 92.75 $\pm$0.17 89.80 $\pm$0.05 92.61 $\pm$0.05
ViT-L/14 95.36 96.13 $\pm$0.06 95.18 $\pm$0.02 95.16 $\pm$0.03
CIFAR-10.1 ViT-B/16 84.00 88.52 $\pm$0.33 83.75 $\pm$0.21 88.72 $\pm$0.33
ViT-L/14 91.20 92.22 $\pm$0.25 91.32 $\pm$0.12 91.02 $\pm$0.02
C-Average ViT-B/16 60.15 68.00 59.75 73.22
ViT-L/14 76.04 79.18 75.01 78.06
Table 10: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with ViT-B/16 and ViT-L/14 as visual encoders.
CLIP TENT TPT (BS=32) CLIPArTT
Original ViT-B/16 64.76 71.73 $\pm$0.14 67.15 $\pm$0.23 71.34 $\pm$0.07
ViT-L/14 73.28 78.03 $\pm$0.08 76.85 $\pm$0.06 79.42 $\pm$0.08
C-Average ViT-B/16 32.01 37.90 33.73 40.08
ViT-L/14 44.59 50.14 47.58 52.52
Table 11: Accuracy (%) on CIFAR-100 and CIFAR-100-C datasets with ViT-B/16 and ViT-L/14 as visual encoders.
CLIP TENT TENT-10 LAME TPT (BS=32) CLIPArTT
3D ViT-B/16 87.04 87.15 $\pm$0.21 88.08 $\pm$0.05 85.83 $\pm$0.06 83.85 $\pm$0.06 88.01 $\pm$0.47
ViT-B/32 84.51 84.75 $\pm$0.17 85.82 $\pm$0.02 83.85 $\pm$0.03 78.45 $\pm$0.05 87.24 $\pm$0.04
ViT-L/14 90.62 90.93 $\pm$0.01 91.33 $\pm$0.01 90.15 $\pm$0.07 91.07 $\pm$0.05 91.18 $\pm$0.01
YT ViT-B/16 86.71 86.96 $\pm$0.02 87.44 $\pm$0.02 86.24 $\pm$0.0 85.69 $\pm$0.05 86.22 $\pm$0.01
ViT-B/32 84.45 84.72 $\pm$0.01 85.46 $\pm$0.02 84.41 $\pm$0.02 82.95 $\pm$0.01 83.46 $\pm$0.03
ViT-L/14 84.93 85.38 $\pm$0.01 86.65 $\pm$0.05 85.08 $\pm$0.04 85.54 $\pm$0.38 85.55 $\pm$0.01
Table 12: Accuracy (%) on the VisDA-C training split (3D) and validation split (YT) with the three different feature extractors.
Performance analysis under simulated and video shifts.  Results on the 3D and YT splits of VisDA-C are shown in Table 12. With respect to pure CLIP, our method achieves an important improvement of accuracy with the three different image encoders. CLIPArTT also achieves a competitive advantage against TENT, LAME and TPT. We observed however that using TENT with 10 iterations (here named TENT-10) is a strong contender, achieving practically the same score on 3D images and outperforming on YouTube frames by a small margin. This can mean that TENT’s entropy minimization can achieve faster improvement with certain domain shifts.
@@ 6 Conclusions
In this study, we introduce CLIPArTT, the first comprehensive Test-Time Adaptation framework designed specifically for Vision-Language Models. By leveraging the model’s predictions as new pseudo-labels, we effectively minimize cross-entropy loss at test-time, thereby enhancing model performance even within a zero-shot setting. To ensure fair comparisons, we also incorporate Test-Time Adaptation models into this new benchmark for VLMs.
A comprehensive ablation study is conducted to determine optimal hyperparameters and gain deeper insights into the various model configurations. Our experimental results demonstrate that CLIPArTT achieves highly competitive performance across TTA datasets, surpassing state-of-the-art approaches. While TENT remains a strong competitor, our model offers enhanced versatility by effectively addressing both natural and severe domain shifts, thus exhibiting robustness across various scenarios.
Exploring the potential of text prompts in classification represents a promising avenue for future research. Investigating alternative methods to fine-tune these text prompts could yield valuable insights. Moreover, extending the study of Test-Time Adaptation to other scenarios, such as segmentation or object detection with Vision-Language Models (VLMs), holds significant promise for advancing our understanding of model adaptability and performance across diverse tasks.
@@
Appendix A Unsupervised clustering
@@
Appendix B Evolution of similarity matrices
A fundamental component of CLIPArTT is the employment of image-to-image and text-to-text similarity matrices to build new pseudo labels. In Fig. 4, we show an example of a batch of five images. While the largest image-to-image similarity values concentrate on the main diagonal, the text-to-text similarity using the $K$ most confident classes for the new text prompt, highly impacts the correlation between images and text (CLIPArTT logits).
Figure 4: Similarity matrices before (a) and after (b) adaptation using $K=5$ classes. There is an increase in the confidence of the logits when using a pseudolabel from the new prompt with the $K$ most confident classes.
Appendix C Comprehensive experimental results

### v2

@@ 5.2 Comparison on different datasets
Natural images.  In the different Tables 7 and 8, CLIPArTT consistently enhances accuracy across CIFAR-10, CIFAR10.1, and CIFAR-100 datasets compared to the baseline (+2%, +3%, +8% respectively with ViT-B/32). This improvement suggests that uncorrupted datasets, which may not have been seen by the model during training, can benefit from adaptation. Consequently, a zero-shot model can leverage adaptation at test-time, relying solely on the model’s predictions. However, while CLIPArTT outperforms the baseline, as well as LAME, TPT and TTDN, it is worth noting that TENT may yield superior results depending on the visual encoder and the number of classes. Entropy minimization used solely on confident results proves to be also effective. Interestingly, the corruptions in which TENT achieves a better performance tend to already have higher accuracy. For a larger dataset such as ImageNet, most methods perform similarly, with the exception of TDA and CALIP, which improve CLIP performance by over 2%. Notably, both TDA and CALIP require hyperparameter tuning, and we used the recommended settings optimized for ImageNet.
Varied styles and textures.  Next we compared methods on two datasets, PACS and OfficeHome, which contain images from very different domains and are often used as benchmark in domain generalization. Once again, these domains do not constitute real distribution shifts for CLIP which was trained with a broad range of datasets. As can be seen, the best performance is achieved by methods such as TTDN for this setting. Nevertheless, it is important to highlight CLIPArTT’s robustness, as it consistently maintains or improves model performance after adaptation. For instance, on the PACS dataset, CLIPArTT achieves an accuracy of 93.95%, compared to 93.65% for the original CLIP.
Backbone CLIP TENT CLIPArTT
CIFAR-10.1 ViT-B/16 84.00 88.52 $\pm$0.33 88.72 $\pm$0.33
ViT-L/14 91.20 92.22 $\pm$0.25 91.02 $\pm$0.02
CIFAR-10-C ViT-B/16 60.15 68.00 $\pm$0.18 73.22 $\pm$0.17
ViT-L/14 76.04 79.18 $\pm$0.10 78.06 $\pm$0.15
CIFAR-100-C ViT-B/16 32.01 37.90 $\pm$0.18 40.08 $\pm$0.18
ViT-L/14 44.59 50.14 $\pm$0.15 52.52 $\pm$0.18
Table 8: Accuracy (%) on CIFAR-10, CIFAR-10.1, CIFAR-10-C, CIFAR-100 and CIFAR-100-C datasets with ViT-B/16 and ViT-L/14 as visual encoders.
@@
Common corruptions.  In the various referenced tables, including Tables 7 and 8, CLIPArTT consistently outperforms the most benchmarks across all visual encoders, resulting in enhancements of up to 13% on average. A more detailed examination reveals that CLIPArTT frequently surpasses TENT, particularly in instances where the baseline performs poorly. This observation implies that CLIPArTT consistently outperforms TENT, which tends to compromise the model’s performance under conditions of low baseline confidence. For instance, in Table 6, TENT experiences a 3% decrease compared to the baseline under Impulse Noise, while CLIPArTT exhibits a 12% improvement. In Figure 3 (right), it is evident that CLIPArTT achieves better performance more rapidly across varying numbers of iterations. Conversely, when the baseline is confident, CLIPArTT consistently maintains a high performance. Finally, our method performs effectively on larger datasets with a greater number of classes, consistently improving upon the baseline by 6% on ImageNet-C.
Simulated images and video.  The performance of CLIPArTT on the YT dataset is suboptimal, a trend that can be observed in scenarios where the CLIP model already exhibits high confidence. However, our observations indicate that on the 3D dataset, CLIPArTT demonstrates a notable competitive advantage over other Test-Time Adaptation (TTA) methods and the baseline, achieving an improvement of nearly 3%. Additionally, while TENT – often considered one of the more robust methods in handling various corruptions – results in a performance decline relative to the baseline, CLIPArTT maintains its effectiveness.
5.3 Limitations
Our results indicate that CLIPArTT performs particularly well under significant domain shifts such as corruptions, though it is slightly less effective on natural images. While the performance of CLIPArTT in these cases is not as strong as that of TENT, it remains stable without deteriorating CLIP’s capabilities, whereas the performance of TENT degrades under severe domain shifts. It is important to mention that any dataset not included in CLIP’s pre-training can be considered a potential domain shift. However, natural images are more likely to have been seen by CLIP during pre-training, making adaptation for them prone to overfitting. This may explain why CLIPArTT excels in more challenging scenarios. To further explore the potential limitations of our method, we have included two additional scenarios in the supplementary material: a) evaluation in a highly imbalanced setting, where only $C$ randomly chosen classes are present in a batch at random, and b) open-set classification, where out-of-distribution images (i.e., from classes not present in the text prompts) are introduced into the batches.
@@ 6 Conclusions
We introduce CLIPArTT, a novel Test-Time Adaptation framework designed specifically for VLMs. By leveraging the model’s predictions as new pseudo-labels, we effectively minimize cross-entropy loss at test-time, thereby enhancing model performance even within a zero-shot setting. To ensure fair comparisons, we also incorporate Test-Time Adaptation models into this new benchmark for VLMs.
A comprehensive ablation study determined optimal hyperparameters and helped gain deeper insights into the various model configurations. Our experimental results demonstrate that CLIPArTT achieves highly competitive performance across TTA datasets, surpassing state-of-the-art approaches. While TENT remains a strong competitor, our model offers enhanced versatility by effectively addressing both natural and severe domain shifts, thus exhibiting robustness across various scenarios.
Exploring the potential of text prompts in classification is a promising avenue for future research. Investigating alternative methods to fine-tune these text prompts could yield valuable insights. Moreover, extending the study of Test-Time Adaptation to other scenarios, such as segmentation or object detection with Vision-Language Models (VLMs), holds significant promise for advancing our understanding of model adaptability and performance across diverse tasks.
@@
Appendix A Dataset Details
We evaluate CLIPArTT’s performance across diverse TTA datasets using established methodologies. These datasets simulate challenging scenarios, offering insights into our approach’s efficacy. Additionally, we explore CLIPArTT’s adaptability on other datasets through zero-shot test-time adaptation.
Our evaluation framework encompasses natural images, varied styles and textures images, common corruptions, simulated images, and video providing a comprehensive assessment of the model’s performance across diverse challenges.
In our evaluation of natural images (also known as zero-shot scenario), we utilize CIFAR-10, CIFAR-10.1, and CIFAR-100 datasets, each comprising 10,000 images and featuring 10 and 100 classes, respectively. These datasets represent natural imagery and are novel to the model under scrutiny. Notably, CIFAR-10.1 introduces a natural domain shift from CIFAR-10, thereby enriching our assessment with varied and nuanced data distributions. We also evaluate our method on ImageNet and extend our investigation on two datasets mostly used in the field of domain generalization: PACS Li et al. (2017) and OfficeHome Venkateswara et al. (2017) datasets, instrumental in understanding texture and style variations. The PACS dataset consists of 9,991 images across four domains (Art, Cartoons, Photos, Sketches) and seven classes. Lastly, the OfficeHome dataset includes 15,588 images across four domains (Art, Clipart, Product, Real) and 65 classes. Evaluating across these distinct scenarios showcases the generalizability of our method.
Transitioning to our investigation of common corruptions, we turn to the CIFAR-10-C and CIFAR-100-C Hendrycks and Dietterich (2019). These datasets offer a diverse range of 15 distinct corruptions, including elastic transform and impulse noise, among others. Each corruption is characterized by 5 severity levels, yielding a total of 75 unique testing scenarios per dataset. Within each severity level, there are 10,000 images, contributing to a comprehensive evaluation of the model’s robustness against a variety of corruption types and intensities. Finally, we test our method on ImageNet-C to evaluate its performance on a larger dataset with 1,000 classes.
Finally, we examine the VisDA-C dataset’s Peng et al. (2018) two domain shifts: simulated (3D) and video (YT). The former comprises a set of 152,397 images rendered in 3D across 12 different classes. The latter includes 72,372 YouTube video frames spanning the same categories. This dataset presents an important challenge, as it bridges the gap of the type of imagery that a model can be applied on.
Appendix B Unsupervised clustering
@@
Appendix C Additional settings
We explore additional experimental settings to further explore the strengths and weaknesses of CLIPArTT. Although these scenarios deviate from the standard practices in the TTA literature, they are useful to evaluate a method’s performance in potentially challenging real-world applications. For most of the following experiments, we utilize the CIFAR-10 dataset’s variants, unless otherwise is mentioned.
C.1 Imbalanced batch instances
CLIPArTT adapts to batch data in a trandsductive manner by computing the image-to-image similarity. The diversity of the batches cannot be ensured due to their finite size. The opposite case naturally arises in large scale classification datasets such as ImageNet, where including at least one image per class in a 128--size batch is impossible. On the other hand, adaptation to highly imbalanced batches is understudied.
In this experiment, we force extreme imbalances in the batches, by allowing only $C$ randomly chosen classes to be present in the batch. As CLIPArTT is dependant also on the most probable predictions, it is expected that the miss-classification due to the imbalance would drive the model to directions opposite to the actual correct classes. In this experiment, we evaluate with $C=\{2,3,4,5\}$ as the possible number of classes to be present in the batches, and compare CLIPArTT against TENT. Final results are shown in Table 9).
It can be observed that TENT shows a higher robustness to class imbalance, as for each prediction, it depends less on the other images’ predictions (i.e., induction). Interestingly, we observe a trend where TENT’s performance decrease when more classes are present (e.g., $C=5$ instead of $C=2$), whereas CLIPArTT increase its performance as $C$ grows.
CLIPArTT TENT
2 3 4 5 2 3 4 5
Original 76.99 81.12 84.37 86.92 97.07 96.14 95.27 94.09
CIFAR-10.1 74.21 78.96 82.57 85.14 96.47 94.33 92.56 90.68
Gaussian Noise 47.12 50.83 54.18 58.63 63.76 49.49 49.92 47.15
Shot Noise 48.67 53.30 56.37 60.24 69.89 55.83 56.68 53.55
Impulse Noise 45.98 49.82 53.15 57.48 66.87 54.47 54.77 52.88
Defocus Blur 60.80 67.78 71.24 74.85 89.19 85.39 83.05 81.44
Glass Blur 47.82 53.07 56.45 60.12 77.11 65.34 62.72 58.78
Motion Blur 60.85 66.82 70.67 73.43 83.33 79.88 79.28 75.89
Zoom Blur 56.78 66.52 70.43 72.94 86.88 84.15 81.89 81.35
Snow 59.20 66.38 71.29 73.97 91.84 87.60 85.32 82.45
Frost 60.02 68.08 72.12 75.15 92.50 88.62 86.18 84.25
Fog 58.57 66.11 70.54 73.12 91.26 87.01 84.07 83.12
Brightness 72.29 77.62 80.75 83.55 95.45 93.82 92.29 91.15
Contrast 62.18 71.68 75.08 77.64 89.75 87.11 85.72 83.25
Elastic Transform 50.85 58.91 62.12 65.12 88.17 82.05 78.13 76.05
Pixelate 51.75 57.20 60.42 63.08 77.35 71.38 69.22 66.85
JPEG Compression 49.63 56.08 59.32 62.05 79.98 72.34 69.85 68.17
Table 9: Accuracy on imbalance datasets with different numbers of available classes at each batch.
C.2 Open-Set classification
Another interesting scenario is open-set classification, where images from classes that are not considered are present in the batch. These images could potentially affect performance, specially in transductive methods and particularly in CLIPArTT, as the image-to-image and text-to-text similarities are combined. The only available text prompts are wrongly assigned to the out-of-distribution (OOD) samples.
To measure the effectiveness of our method in this scenario, we utilize SVHN-C, a corrupted version of the SVHN Netzer et al. (2011) dataset, analogous to CIFAR-10.C. For each batch of 128 image, 128 additional OOD images are included. Both parts are used at the same time for adaptation, and accuracy is only measured on the first half.
As seen in Table 10, CLIPArTT defeats TENT’s open-set accuracy by a significant margin. This encouraging result suggests that our method is robust to semantic out-of-distribution perturbations.
CLIPArTT TENT
Gaussian Noise 59.72 $\pm$0.05 41.27 $\pm$0.03
Shot Noise 62.11 $\pm$0.07 48.14 $\pm$0.05
Impulse Noise 55.23 $\pm$0.10 47.86 $\pm$0.07
Defocus Blur 75.44 $\pm$0.07 72.03 $\pm$0.05
Glass Blur 60.12 $\pm$0.07 42.08 $\pm$0.08
Motion Blur 74.60 $\pm$0.05 65.71 $\pm$0.05
Zoom Blur 76.15 $\pm$0.05 71.45 $\pm$0.06
Snow 74.69 $\pm$0.05 73.99 $\pm$0.08
Frost 77.72 $\pm$0.05 74.49 $\pm$0.09
Fog 72.87 $\pm$0.03 70.57 $\pm$0.04
Brightness 85.70 $\pm$0.09 82.34 $\pm$0.05
Contrast 76.49 $\pm$0.04 71.67 $\pm$0.04
Elastic Transform 67.42 $\pm$0.01 65.99 $\pm$0.12
Pixelate 62.02 $\pm$0.11 54.25 $\pm$0.08
JPEG Compression 60.81 $\pm$0.07 54.09 $\pm$0.10
Average 69.41 $\pm$8.86 62.40 $\pm$13.20
Table 10: Open-set accuracy on CIFAR-10-C when using SVHN-C as the OOD dataset.
C.3 Generalization after adaptation
In TTA, the classification performance is evaluated directly on the adapted set of images in an episodic manner. However, a high accuracy on this set does not necessarily guarantee a good performance in a separate set of images unknown to the model.
[…]
While TENT obtains a higher accuracy on the adaptation splits of natural images (i.e., CIFAR-10 and CIFAR-10.1), following a similar trend as in the main experiments’ results, the difference in the generalization accuracy is smaller with respect to CLIPArTT’s. Moreover, our method demonstrates a consistently better generalization and adaptation accuracy on corrupted samples, representing stronger domain shifts.
CLIPArTT TENT
Acc. Acc. Gen. Acc. Acc. Gen.
Original 95.43 $\pm$0.01 71.23 $\pm$0.0008 97.02 $\pm$0.02 72.68 $\pm$0.0023
CIFAR-10.1 88.75 $\pm$0.85 65.87 $\pm$0.8083 90.49 $\pm$0.38 67.07 $\pm$1.1
Gaussian Noise 63.79 $\pm$0.20 46.87 $\pm$0.0008 45.30 $\pm$0.05 30.83 $\pm$0.0030
Shot Noise 66.38 $\pm$0.10 79.80 $\pm$0.0038 50.65 $\pm$0.18 35.85 $\pm$0.0025
Impulse Noise 59.31 $\pm$0.19 44.56 $\pm$0.0023 51.91 $\pm$0.20 37.99 $\pm$0.0037
Defocus Blur 81.26 $\pm$0.03 60.13 $\pm$0.0016 84.48 $\pm$0.07 60.32 $\pm$0.0028
Glass Blur 65.38 $\pm$0.04 49.47 $\pm$0.0020 56.68 $\pm$0.26 41.32 $\pm$0.0035
Motion Blur 80.41 $\pm$0.12 60.08 $\pm$0.0022 75.70 $\pm$0.13 55.72 $\pm$0.0013
Zoom Blur 81.97 $\pm$0.05 60.63 $\pm$0.0009 81.14 $\pm$0.13 59.33 $\pm$0.0007
Snow 82.06 $\pm$0.14 60.83 $\pm$0.0014 83.08 $\pm$0.13 61.08 $\pm$0.0008
Frost 83.85 $\pm$0.19 61.73 $\pm$0.0009 84.70 $\pm$0.11 66.75 $\pm$0.0006
Fog 90.04 $\pm$0.27 59.48 $\pm$0.0027 81.56 $\pm$0.09 60.69 $\pm$0.0033
Brightness 91.73 $\pm$0.04 68.01 $\pm$0.0017 92.93 $\pm$0.03 68.80 $\pm$0.0018
Contrast 82.63 $\pm$0.07 61.28 $\pm$0.0025 83.53 $\pm$0.06 62.97 $\pm$0.0009
Elastic Transform 74.04 $\pm$0.22 54.87 $\pm$0.0011 74.08 $\pm$0.10 54.92 $\pm$0.0009
Pixelate 70.58 $\pm$0.02 51.53 $\pm$0.0042 67.35 $\pm$0.18 50.19 $\pm$0.0004
JPEG Compression 67.32 $\pm$0.06 49.56 $\pm$0.0023 66.25 $\pm$0.14 48.85 $\pm$0.0021
Average 75.38$\pm$9.04 55.92 $\pm$6.62 71.78 $\pm$14.14 52.77 $\pm$11.10
Table 11: Generalization results on CIFAR-10 variants. Each batch is divided into an adaptation split and a validation split. Accuracy (Acc.) is measured on the former after adaptation, whilst the generalization accuracy (Acc. Gen.) is measured on the later.
Appendix D Computational Cost
In this section, we compare the computational cost of CLIPArTT with other TTA methods through a thorough evaluation under consistent conditions, using an NVIDIA A6000 GPU within the same Python environment. The provided table 12 compares adaptation time, memory usage, and the number of learnable parameters across various TTA methods, including our proposed CLIPArTT. The results demonstrate that CLIPArTT maintains competitive adaptation time and memory usage relative to other approaches, such as TENT and TPT.
Method Adaptation Time Memory Pct. of Learnable Parameters
TENT 0.28 s 1.5 GB 0.026%
TPT 0.26 s 1.7 GB 0.001%
CLIPArTT 0.55 s 1.7 GB 0.026%
Table 12: Comparison of Computational Cost.
Appendix E Comprehensive experimental results